the temperature of a gas started at 312 with a yolume of 4.5 L increased to 372 with a new volume of 1.2 what would the new pressure be if the initial pressure was 1.1 atm


1
Expert's answer
2021-05-26T07:07:20-0400

"\frac {P_1V_1} {T_1} =\frac {P_2 V_2}{T_2}"


"P_2=\frac {1.1×4.5×372}{312×1.2}"


"P_2 =4.9atm"